Reeves Company is preparing its master budget for July. Use the given estimates to determine the amounts necessary for each of the following requirements. (Estimates may be related to more than one requirement.)

a. What should total sales revenue be if territories A and B estimate sales of 8,000 and 20,000 units, respectively, and the unit selling price is $55?

$Answer

b. If the beginning finished goods inventory is an estimated 2,500 units and the desired ending inventory is 1,000 units, how many units should be produced?

Answer

c. What dollar amount of material should be purchased at $3 per pound if each unit of product requires 2 pounds and beginning and ending materials inventories should be 3,000 and 4,000 pounds, respectively?

$Answer

d. How much direct labor cost should be incurred if each unit produced requires 1.5 hours at an hourly rate of $16?

$Answer

e. How much manufacturing overhead should be incurred if fixed manufacturing overhead is $60,000 and variable manufacturing overhead is $2 per direct labor hour?

$Answer

Save AnswersNext
Answer #1
Units Selling Price Revenue
a) A 8000 55           4,40,000
B 20000 55        11,00,000
       15,40,000
b) Beginning Inventory 2500
Closing inventory 1000
Sales (8000+20000) 28000
Sales = (Opening+Production-closing)
28000 = 2500+Production- 1000
28000-1500 = Production
         26,500 = Production
c) Beginning material invenory = 3000
Closing material invenory = 4000
Each unit requirement of material = 2
Requirement of 26500 units (26500*2) = 53000
Purchase of material = 53000-3000+4000
= 54000 units
rate per pound = 3
Cost of purchase of material           1,62,000
d) Direct labour cost = 26500*1.5*16
=           6,36,000
e) Fixed manufacturing overhead 60000
variable manufacturing overhead 2*26500*1.5
79500
Total 60000+79500
          1,39,500